OpenStudy (anonymous):

can someone showme how to simplify (sqrt 10/3)/(sqrt 5)

OpenStudy (anonymous):

hey i got expelled to

OpenStudy (anonymous):

\[\frac{\sqrt{10/3}}{\sqrt{5}}\] \[=\sqrt{(\frac{10/3}{5})}\] \[= \sqrt{10/15} = \sqrt{2/3}\]

OpenStudy (anonymous):

I bet you did...

OpenStudy (anonymous):

so

OpenStudy (anonymous):

i got expelled cuz i touched ur mom

OpenStudy (anonymous):

Thanks polpak. gordita you better crunch your fat Azz back to mexico.

OpenStudy (anonymous):

thats not what your mom said